# Hierarchical Decision Mamba

### v1

@@ 5.1 Time Comparison
Mamba models have outpaced Transformers in terms of speed in other applications. Also, HDT and HDM require the training of two models, and the extra computational cost may not be worth the performance benefits. Because of this, we compare the time required to perform a training iteration and the inference step using the different methods across the 7 task environments available in the D4RL benchmark. We use a batch size of 16, an embedding size of 128, a sequence length of 20, and 6 layers per model. Specifically, for the HDT and the HDM, we employ 6 layers for both the high-level and low-level models, and we measure the time to train both models. For training, we measure the time it takes for a gradient calculation and update. For inference, we measure the time it takes to build the sequences, obtain an action from the model and perform the transition. To ensure statistical robustness, we repeat both these steps 1000 times for each model, presenting the average and standard deviation time to perform a training iteration, and an inference step in Table 2. Results show that during training there’s not much difference between the Mamba methods and their Transformer predecessors. As expected the HDT and the HDM take close to double the time to train due to having double the models than the DT and the DM, respectively. Adding rewards to the DM does not increase the training time significantly. This concludes that one should use the DM with RTG as the performance increase compensates for the small computational cost. At inference time however, the Mamba methods are faster than the Transformer methods. In addition to the increase in performance, this computational boost further shows the benefits of our methods compared to the baselines.
Table 2: Average and STD time for a single training iteration, and to perform inference of an episode, across the different D4RL tasks, using a batch size of 16, of each of the 5 methods, configured with 6 layers, an embedding size of 128 and sequence length 20.
@@ 6 Conclusion
In summary, we introduced the DM and the HDM models, evolutions of the existing state-of-the-art sequence models, DT and HDT, respectively, by leveraging the potent Mamba architecture. We show that the DM does not rely on the sequence of RTG to perform, although including it does slightly increases its performance. Through our evaluation across seven diverse tasks within the D4RL benchmark and varying the demonstration data set, we demonstrate the superiority of these Mamba models over their Transformer-based predecessors in the majority of cases. Specifically, DM and HDM outperform their transformer counterparts in most settings, while also being simpler and faster to train and perform inference. Lastly, HDM outperforms the baselines in most tasks, while being task-independent by not requiring user specified values. This advancement underscores the viability of Mamba architectures in behaviour cloning sequence modelling, and opens the way to keep solving more complex sequence modelling problems through imitation learning.
@@ Acknowledgments
This work is supported by NOVA LINCS (UIDB/04516/2020) with the financial support of ’FCT - Fundação para a Ciência e Tecnologia’ and also through the research grant ’2022.14197.BD’.

### v2

@@ 5.1 Time Comparison
Mamba models have outpaced Transformers in terms of speed in other applications. Also, HDT and HDM require the training of two models, and the extra computational cost may not be worth the performance benefits. Because of this, we compare the time required to perform a training iteration and the inference step using the different methods across the 7 task environments available in the D4RL benchmark. We use a batch size of 16, an embedding size of 128, a sequence length of 20, and 6 layers per model. Specifically, for the HDT and the HDM, we employ 6 layers for both the high-level and low-level models, and we measure the time to train both models. For training, we measure the time it takes for a gradient calculation and update. For inference, we measure the time it takes to build the sequences, obtain an action from the model and perform the transition. To ensure statistical robustness, we repeat both these steps 1000 times for each model, presenting the average and standard deviation time to perform a training iteration, and an inference step in Table 2. Results show that during training there’s not much difference between the Mamba methods and their Transformer predecessors. As expected the HDT and the HDM take close to double the time to train due to having double the models than the DT and the DM, respectively. Adding rewards to the DM does not increase the training time significantly. At inference time however, the Mamba methods are faster than the Transformer methods. In addition to the increase in performance, this computational boost further shows the benefits of our methods compared to the baselines.
Table 2: Average and STD time for a single training iteration, and to perform inference of an episode, across the different D4RL tasks, using a batch size of 16, of each of the 5 methods, configured with 6 layers, an embedding size of 128 and sequence length 20. Lowest values are highlighted in bold.
@@ 6 Conclusion
In summary, we introduced the DM and the HDM models, evolutions of the existing state-of-the-art sequence models, DT and HDT, respectively, by leveraging the potent Mamba architecture. We show that the DM does not rely on the sequence of RTG, or a reward function to perform. Through our evaluation across seven diverse tasks within the D4RL benchmark and varying the demonstration data set, we demonstrate the superiority of these Mamba models over their Transformer-based predecessors in the majority of cases. Specifically, DM and HDM outperform their transformer counterparts in most settings, while also being faster to train and perform inference. Lastly, DM outperforms the baselines in most tasks, while being task-independent by not requiring user specified values. This advancement underscores the viability of Mamba architectures in behaviour cloning sequence modelling, and opens the way to keep solving more complex sequence modelling problems through imitation learning.
@@ Acknowledgments
This work is supported by NOVA LINCS ref. UIDB/04516/2020
(https://doi.org/10.54499/UIDB/04516/2020) and ref. UIDP/04516/2020 (https://doi.org/10.54499/UIDP/04516/2020) with the financial support of FCT.IP, and also through the research grant 2022.14197.BD.